An elephant lifts a fallen tree by exerting a force of 8.25 x 103​ ​ N upward. If she lifts the tree 2.10 m in 1.05 seconds, how much power does she use?
1
Expert's answer
2020-11-19T06:01:39-0500

By the definition of the power, we have:


"P=\dfrac{W}{t}=\dfrac{Fh}{t},""P=\dfrac{8.25\cdot 10^3\ N\cdot 2.10\ m}{1.05\ s}=1.65\cdot 10^4\ W."

Answer:

"P=1.65\cdot 10^4\ W."